# Effect of Off-diagonal NSI Parameters on Entanglement Measurements in Neutrino Oscillations

### v1

@@ V.1 Analysis in probability level
Now we analyze the impact of the NSI parameter $\epsilon_{e\tau}$ on the appearance probability. The middle panel of Fig. 1 illustrates this effect, where the upper row corresponds to $\delta_{CP}=212^{\circ}$ and the lower row to $\delta_{CP}=0^{\circ}$. The red curve represents the standard oscillation scenario, the green curve shows the probability in the presence of non-zero $\epsilon_{e\tau}$, and the blue curve corresponds to the case with $\epsilon_{e\tau}=0.27$ and $\phi_{e\tau}=1.62\pi$. For $\delta_{CP}=212^{\circ}$, the green curve has a significantly lower amplitude than the red curve, indicating a suppression in the appearance probability due to the presence of $\epsilon_{e\tau}$. The blue curve, on the other hand, closely follows the red curve, suggesting only a minor effect from the inclusion of the NSI phase in this case.
However, when $\delta_{CP}=0^{\circ}$ in Fig. 1, the overall pattern changes. The green curve now shows a much higher amplitude than the red curve, while the blue curve exhibits the lowest amplitude among the three. This change highlights the sensitivity of the appearance probability to both the magnitude and phase of the NSI parameter $\epsilon_{e\tau}$, as well as its strong interplay with the CP-violating phase $\delta_{CP}$.
@@
To understand the behavior of the effect of $\epsilon_{e\tau}$ on $P_{\mu e}$, we split each term of the appearance probability given in eq. V.1 as, by taking $\epsilon_{e\tau}$ and $\phi_{e\tau}$ only, keeping $\epsilon_{e\mu}=0$,
$P(\nu_{\mu}\to\nu_{e})|_{\epsilon_{e\mu,\phi_{e\mu}=0}}$ $=SO+D11+E11+F11+G11+J11+K11+L11,$ (20)
@@
$D11=4\hat{A}\epsilon_{e\tau}s_{23}c_{23}xf^{2}\cos(\phi_{e\tau}+\delta_{CP}),~{}~{}~{}~{}E11=-4\hat{A}\epsilon_{e\tau}s_{23}c_{23}xfg\cos(\Delta+\phi_{e\tau}+\delta_{CP}),$
$F11=-4\hat{A}\epsilon_{e\tau}s_{23}c_{23}yg^{2}\cos\phi_{e\tau},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}G11=4\hat{A}\epsilon_{e\tau}s_{23}c_{23}ygf\cos(\Delta-\phi_{e\tau}),$
$J11=4\hat{A}^{2}g^{2}\epsilon_{e\tau}^{2}s^{2}_{23}c^{2}_{23},~{}~{}~{}~{}K11=4\hat{A}^{2}f^{2}\epsilon_{e\tau}^{2}s^{2}_{23}c^{2}_{23},~{}~{}~{}~{}L11=-8\hat{A}^{2}fg\epsilon_{e\tau}^{2}s^{2}_{23}c^{2}_{23}\cos\Delta.$ (21)
@@
Figure 3 presents the energy dependence of the individual terms contributing to the appearance probability. The upper and lower panels correspond to $\delta_{CP}=212^{\circ}$ and $\delta_{CP}=0^{\circ}$, respectively, while the left and right columns show results for $\phi_{e\tau}=0$ and $\phi_{e\tau}=1.62\pi$. In the case of $\delta_{CP}=212^{\circ}$, it is evident that the terms D11, E11, and L11 dominate the behavior of $P_{\mu e}$, with the remaining terms contributing negligibly. For $\phi_{e\tau}=0$, all three dominant terms, D11, E11, and L11 exhibit negative amplitudes throughout the energy range, leading to an overall suppression of the appearance probability compared to the standard oscillation scenario. This observation aligns with the green curve in the middle panel of the upper row in Fig. 1.
When $\phi_{e\tau}=1.62\pi$, the E11 term becomes positive, while D11 and L11 remain negative. This results in a slight enhancement of the overall amplitude compared to the case with $\phi_{e\tau}=0$, although the appearance probability is still close to that of standard oscillations, as shown by the blue curve in the corresponding panel of Fig. 1.
@@
For the CP-conserving case with $\delta_{CP}=0^{\circ}$, the behavior changes considerably. The lower panel of Fig. 3 reveals that when $\phi_{e\tau}=0$, D11 and E11 contribute positively, while L11 remains negative. The combined effect of these terms leads to a significantly higher appearance probability than that of standard oscillations, which is consistent with the green curve in the middle column of lower row of Fig. 1. However, for $\phi_{e\tau}=1.62\pi$, E11 and L11 contribute large negative amplitudes, while D11 shows positive amplitude. This results in a strong suppression of the overall appearance probability, reflected in the blue curve in the corresponding panel of Fig. 1.
@@
Finally, we examine the influence of $\epsilon_{\mu\tau}$ on the appearance probability. The right-most panels of the upper and lower rows of Fig. 1 illustrate this effect for $\delta_{CP}=212^{\circ}$ and $\delta_{CP}=0^{\circ}$, respectively. The color scheme remains consistent with the previous panels: the red curve corresponds to standard oscillations, the green curve represents the scenario with non-zero $\epsilon_{\mu\tau}$ and $\phi_{\mu\tau}=0$, and the blue curve shows the result for $\epsilon_{\mu\tau}=0.35$ and $\phi_{\mu\tau}=0.6\pi$. As observed in both CP-violating and CP-conserving cases, the green and blue curves exhibit only slight deviations from the red curve. This indicates that the presence of $\epsilon_{\ mu\tau}$ leads to relatively minor modifications in the appearance probability. This behavior is consistent with the analytical expression of $P_{\mu e}$ given in Eq. V.1, which shows that $\epsilon_{\mu\tau}$ does not contribute to the probability up to $\mathcal{O}(\epsilon^{2})$. However, if the expression is expanded to higher orders, specifically beyond $\mathcal{O}(\epsilon^{2})$, subleading terms involving $\epsilon_{\ mu\tau}$ may appear. These higher-order contributions can account for the small deviations observed in the rightmost column of Fig. 1.
@@
Now, we explore the impact of NSI parameters on the disappearance probability. According to Eq. V.1, the disappearance channel $P_{\mu\mu}$ is primarily influenced by $\epsilon_{\mu\tau}$ and its associated phase $\phi_{\mu\tau}$. Figure 4 illustrates the effects of three different NSI parameters on the disappearance probability separately. The left and middle columns show the effects of $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$, respectively, while the rightmost column presents the impact of $\epsilon_{\mu\tau}$. The color scheme remains consistent throughout: the red curve represents the standard oscillation scenario, the green curve shows the result when $\epsilon_{\alpha\beta}$ is non-zero (with the corresponding phase set to zero), and the blue curve corresponds to the case where both $\epsilon_{\alpha\beta}$ and $\phi_{\alpha\beta}$ are non-zero. The shaded area represents the $\nu_{\mu}$ flux for the DUNE experiment in arbitrary units.
@@
From the figure (Fig. 4), it is evident that when either $\epsilon_{e\mu}$ or $\epsilon_{e\tau}$ is present, the green and blue curves show negligible deviations from the standard oscillation (red curve). This observation is consistent with the analytical form of Eq. V.1, which shows independent behaviour on these parameters in the disappearance channel. However, the presence of $\epsilon_{\mu\tau}$ and its phase leads to a noticeable change in the disappearance probability. The rightmost column clearly shows a significant deviation of the green curve from the red curve, indicating the strong effect of $\epsilon_{\mu\tau}$ when the phase is set to zero. Interestingly, the blue curve, which includes both $\epsilon_{\mu\tau}$ and $\phi_{\mu\tau}$, mostly overlaps with the red curve. Furthermore, there is an observable shift in the oscillation minimum for the green curve relative to the red and blue curves, highlighting the sensitivity of $P_{\mu\mu}$ to $\epsilon_{\mu\tau}$ in the absence of a complex phase. To understand the nature, we need to split the modified expression of $P_{\mu\mu}$,
@@
$P1=-2\hat{A}\epsilon_{\mu\tau}\cos\phi_{\mu\tau}\Big{[}\sin^{3}2\theta_{23}\Delta\sin 2\Delta+2\sin 2\theta_{23}\cos^{2}2\theta_{23}\sin^{2}\Delta\Big{]}$
$P2=-2\hat{A}^{2}\sin^{2}2\theta_{23}\epsilon_{\mu\tau}^{2}\Big{[}2\sin^{2}2\theta_{23}\cos^{2}\phi_{\mu\tau}\Delta^{2}\cos 2\Delta+\sin^{2}\phi_{\mu\tau}\Delta\sin 2\Delta\Big{]}.$ (23)
@@
Figure 5 illustrates the contribution of each term to the disappearance probability. Let us first consider the scenario where $\epsilon_{\mu\tau}=0.35$ and $\phi_{\mu\tau}=0$. In this case, the left panel of Fig. 5 shows that the term P1 causes a noticeable shift in the oscillation minimum compared to the standard oscillation. Additionally, beyond the oscillation minimum, the P1 curve exhibits a negative amplitude, leading to an overall suppression of the probability relative to the standard case.
@@
However, when the phase is set to $\phi_{\ mu\tau}=0.6\pi$, both P1 and P2 have a negligible impact on the standard oscillation behavior. As a result, the corresponding blue curve in the rightmost column of Fig. 4 nearly overlaps with the standard oscillation curve, indicating minimal deviation.
V.2 Analysis based on quantum entanglement measures
In this subsection, we discuss how the off-diagonal NSI parameters influence the entanglement measures: Entanglement of Formation (EOF), Concurrence, and Negativity. Since all three entanglement measures are functions of neutrino oscillation and survival probabilities, the impact of NSI parameters on these quantities is essentially determined by their influence on the oscillation probabilities.

### v2

@@ V.1 Analysis in probability level
Now we analyze the impact of the NSI parameter $\epsilon_{e\tau}$ on the appearance probability. The middle panel of Fig. 1 illustrates this effect, where the upper row corresponds to $\delta_{CP}=212^{\circ}$ and the lower row to $\delta_{CP}=0^{\circ}$. The red curve represents the standard oscillation scenario, the green curve shows the probability in the presence of non-zero $\left|\epsilon_{e\tau}\right|$, and the blue curve corresponds to the case with $\left|\epsilon_{e\tau}\right|$ = 0.27 and $\phi_{e\tau}=1.62\pi$. For $\delta_{CP}=212^{\circ}$, the green curve has a significantly lower amplitude than the red curve, indicating a suppression in the appearance probability due to the presence of $\epsilon_{e\tau}$. The blue curve, on the other hand, closely follows the red curve, suggesting only a minor effect from the inclusion of the NSI phase in this case. However, when $\delta_{CP}=0^{\circ}$ in Fig. 1, the overall pattern changes. The green curve now shows a much higher amplitude than the red curve, while the blue curve exhibits the lowest amplitude among the three. This change highlights the sensitivity of the appearance probability to both the magnitude and phase of the NSI parameter $\epsilon_{e\tau}$, as well as its strong interplay with the CP-violating phase $\delta_{CP}$.
@@
To understand the behavior of the effect of $\left|\epsilon_{e\tau}\right|$ on $P_{\mu e}$, we split each term of the appearance probability given in Eq. V.1 as, by taking $\epsilon_{e\tau}$ and $\phi_{e\tau}$ only, keeping $\epsilon_{e\mu}=0$,
$P(\nu_{\mu}\to\nu_{e})|_{\left|\epsilon_{e\mu}\right|,\phi_{e\mu}=0}$ $=SO+D11+E11+F11+G11+J11+K11+L11,$ (26)
@@
$D11=4\hat{A}\left|\epsilon_{e\tau}\right|s_{23}c_{23}xf^{2}\cos(\phi_{e\tau}+\delta_{CP}),~~~~E11=-4\hat{A}\left|\epsilon_{e\tau}\right|s_{23}c_{23}xfg\cos(\Delta+\phi_{e\tau}+\delta_{CP}),$
$F11=-4\hat{A}\left|\epsilon_{e\tau}\right|s_{23}c_{23}yg^{2}\cos\phi_{e\tau},~~~~~~~~~~~~G11=4\hat{A}\left|\epsilon_{e\tau}\right|s_{23}c_{23}ygf\cos(\Delta-\phi_{e\tau}),$
$J11=4\hat{A}^{2}g^{2}\left|\epsilon_{e\tau}\right|^{2}s^{2}_{23}c^{2}_{23},~~~~K11=4\hat{A}^{2}f^{2}\left|\epsilon_{e\tau}\right|^{2}s^{2}_{23}c^{2}_{23},~~~~L11=-8\hat{A}^{2}fg\left|\epsilon_{e\tau}\right|^{2}s^{2}_{23}c^{2}_{23}\cos\Delta.$
@@
Figure 3 presents the energy dependence of the individual terms contributing to the appearance probability. The upper and lower panels correspond to $\delta_{CP}=212^{\circ}$ and $\delta_{CP}=0^{\circ}$, respectively, while the left and right columns show results for $\phi_{e\tau}=0$ and $\phi_{e\tau}=1.62\pi$. In the case of $\delta_{CP}=212^{\circ}$, it is evident that the terms D11, E11, and L11 dominate the behavior of $P_{\mu e}$, with the remaining terms contributing negligibly. For $\phi_{e\tau}=0$, all three dominant terms, D11, E11, and L11 exhibit negative amplitudes throughout the energy range, leading to an overall suppression of the appearance probability compared to the standard oscillation scenario. This observation aligns with the green curve in the middle panel of the upper row in Fig. 1. When $\phi_{e\tau}=1.62\pi$, the E11 term becomes positive, while D11 and L11 remain negative. This results in a slight enhancement of the overall amplitude compared to the case with $\phi_{e\tau}=0$, although the appearance probability is still close to that of standard oscillations, as shown by the blue curve in the corresponding panel of Fig. 1.
@@
For the CP-conserving case with $\delta_{CP}=0^{\circ}$, the behavior changes considerably. The lower panel of Fig. 3 reveals that when $\phi_{e\tau}=0$, D11 and E11 contribute positively, while L11 remains negative. The combined effect of these terms leads to a significantly higher appearance probability than that of standard oscillations, which is consistent with the green curve in the middle column of the lower row of Fig. 1. However, for $\phi_{e\tau}=1.62\pi$, E11 and L11 contribute large negative amplitudes, while D11 shows positive amplitude. This results in a strong suppression of the overall appearance probability, reflected in the blue curve in the corresponding panel of Fig. 1.
@@
Finally, we examine the influence of $\epsilon_{\mu\tau}$ on the appearance probability. The right-most panels of the upper and lower rows of Fig. 1 illustrate this effect for $\delta_{CP}=212^{\circ}$ and $\delta_{CP}=0^{\circ}$, respectively. The color scheme remains consistent with the previous panels: the red curve corresponds to standard oscillations, the green curve represents the scenario with non-zero $\left|\epsilon_{\mu\tau}\right|$ and $\phi_{\mu\tau}=0$, and the blue curve shows the result for $\left|\epsilon_{\mu\tau}\right|$ = 0.35 and $\phi_{\mu\tau}=0.6\pi$. As observed in both CP-violating and CP-conserving cases, the green and blue curves exhibit only slight deviations from the red curve. This indicates that the presence of $\epsilon_{\mu\tau}$ leads to relatively minor modifications in the appearance probability. This behavior is consistent with the analytical expression of $P_{\mu e}$ given in Eq. V.1, which shows that $\epsilon_{\mu\tau}$ does not contribute to the probability up to $\mathcal{O}(\epsilon^{2})$. However, if the expression is expanded to higher orders, specifically beyond $\mathcal{O}(\epsilon^{2})$, subleading terms involving $\epsilon_{\mu\tau}$ may appear. These higher-order contributions can account for the small deviations observed in the rightmost column of Fig. 1.
@@
Now, we explore the impact of NSI parameters on the disappearance probability. According to Eq. V.1, the disappearance channel $P_{\mu\mu}$ is primarily influenced by $\left|\epsilon_{\mu\tau}\right|$ and its associated phase $\phi_{\mu\tau}$. Figure 4 illustrates the effects of three different NSI parameters on the disappearance probability separately. The left and middle columns show the effects of $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$, respectively, while the rightmost column presents the impact of $\epsilon_{\mu\tau}$. The color scheme remains consistent throughout: the red curve represents the standard oscillation scenario, the green curve shows the result when $\left|\epsilon_{\alpha\beta}\right|$ is non-zero with the corresponding phase set to zero, and the blue curve corresponds to the case where both $\left|\epsilon_{\alpha\beta}\right|$ and $\phi_{\alpha\beta}$ are non-zero. The shaded area represents the $\nu_{\mu}$ flux for the DUNE experiment in arbitrary units.
Figure 4: The disappearance probability as a function of neutrino energy for the DUNE experiment with $\delta_{CP}=212^{\circ}$. Left (middle) [right] panel represent the curves with one off-diagonal parameter one at a time, $\epsilon_{e\mu}$ ($\epsilon_{e\tau}$) $\epsilon_{\mu\tau}$] and their associated phases. Color codes are given in the legend.
Figure 5: The effect of each term containing $\left|\epsilon_{\mu\tau}\right|$ and $\phi_{\mu\tau}$ in disappearance probability as a function of neutrino energy with $\delta_{CP}=212^{\circ}$. Left (right) panel shows the results with $\phi_{\mu\tau}=0~(0.6\pi$ rad). Color codes are given in the legend. For more about each term, see the text.
@@
From the figure (Fig. 4), it is evident that when either $\epsilon_{e\mu}$ or $\epsilon_{e\tau}$ is present, the green and blue curves show negligible deviations from the standard oscillation (red curve). This observation is consistent with the analytical form of Eq. V.1, which shows independent behaviour on these parameters in the disappearance channel. However, the presence of $\left|\epsilon_{\mu\tau}\right|$ and its phase leads to a noticeable change in the disappearance probability. The rightmost column clearly shows a significant deviation of the green curve from the red curve, indicating the strong effect of $\left|\epsilon_{\mu\tau}\right|$ when the phase is set to zero. Interestingly, the blue curve, which includes both $\left|\epsilon_{\mu\tau}\right|$ and $\phi_{\mu\tau}$, mostly overlaps with the red curve. Furthermore, there is an observable shift in the oscillation minimum for the green curve relative to the red and blue curves, highlighting the behavior of $P_{\mu\mu}$ to $\left|\epsilon_{\mu\tau}\right|$ in the absence of a complex phase. To understand the nature, we need to split the modified expression of $P_{\mu\mu}$,
@@
$P1=-2\hat{A}\left|\epsilon_{\mu\tau}\right|\cos\phi_{\mu\tau}\Big[\sin^{3}2\theta_{23}\Delta\sin 2\Delta+2\sin 2\theta_{23}\cos^{2}2\theta_{23}\sin^{2}\Delta\Big]$
$P2=-2\hat{A}^{2}\sin^{2}2\theta_{23}\left|\epsilon_{\mu\tau}\right|^{2}\Big[2\sin^{2}2\theta_{23}\cos^{2}\phi_{\mu\tau}\Delta^{2}\cos 2\Delta+\sin^{2}\phi_{\mu\tau}\Delta\sin 2\Delta\Big].$ (29)
@@
Figure 5 illustrates the contribution of each term to the disappearance probability. Let us first consider the scenario where $\left|\epsilon_{\mu\tau}\right|$ = 0.35 and $\phi_{\mu\tau}=0$. In this case, the left panel of Fig. 5 shows that the term P1 causes a noticeable shift in the oscillation minimum compared to the standard oscillation. Additionally, beyond the oscillation minimum, the P1 curve exhibits a negative amplitude, leading to an overall suppression of the probability relative to the standard case.
@@
However, when the phase is set to $\phi_{\mu\tau}=0.6\pi$, both P1 and P2 have a negligible impact on the standard oscillation behavior. As a result, the corresponding blue curve in the rightmost column of Fig. 4 nearly overlaps with the standard oscillation curve, indicating minimal deviation.
V.2 Analysis in event level
To further complement the probability level study, we simulate the expected event rates at DUNE using the General Long-Baseline Experiment Simulator (GLoBES) framework Huber et al. (2005, 2007), where the probability engine has been modified to include the effects of off-diagonal NSI parameters. While probabilities provide the underlying theoretical description of oscillations, experiments measure the corresponding event rates. However, these two quantities are not independent, the event rates are derived from oscillation probabilities by incorporating the source flux, neutrino interaction cross sections, and detector response. Consequently, the trends observed at the probability level are directly reflected in the event rates.
Figure 6: Figure shows the event rates for appearance (upper) and disappearance (lower) channels as a function of neutrino energy for the DUNE experiment with $\delta_{CP}=212^{\circ}$. Left (middle) [right] panel represent the curves with one off-diagonal parameter at a time, $\epsilon_{e\mu}$ ($\epsilon_{e\tau}$) $\epsilon_{\mu\tau}$] and their associated phases. Color codes are given in the legend.
The general expression for the differential event rate in a given interaction channel can be written as Huber et al. (2005, 2007)
$\frac{dn^{\mathrm{IT}}_{\beta}}{dE^{\prime}}=N\int_{0}^{\infty}dE\int_{0}^{\infty}d\hat{E}\,\underbrace{\Phi_{\alpha}(E)}_{\text{Source}}\,\underbrace{\frac{1}{L^{2}}P(\alpha\rightarrow\beta)(E,L,\rho;\theta_{12},\theta_{13},\theta_{23},\Delta m^{2}_{31},\Delta m^{2}_{21},\delta_{\mathrm{CP}})}_{\text{Propagation}},$ (30)
$\times\underbrace{\sigma^{\mathrm{IT}}_{f}(E)\,k^{\mathrm{IT}}_{f}(E-\hat{E})}_{\text{Interaction}}\times\underbrace{T_{f}(\hat{E})\,V_{f}(\hat{E}-E^{\prime})}_{\text{Detection}}.$
Here, $\alpha$ denotes the initial neutrino flavor and $\beta$ the detected flavor. The quantity $\Phi_{\alpha}(E)$ corresponds to the neutrino flux at the source of the initial flavor $\alpha$, $L$ is the baseline length, $N$ represents an overall normalization factor, and $\rho$ is the matter density along the path. The different energies in this formula have different meanings, such as $E$ is the true energy of the incoming neutrino, which cannot be directly observed, $\hat{E}$ is the energy of the produced secondary particle and $E^{\prime}$ denotes the reconstructed neutrino energy, i.e., the experimentally measured value.
Fig. 6 displays the simulated event rates at DUNE as functions of neutrino energy, considering the influence of the off-diagonal NSI parameters $\epsilon_{e\mu}$, $\epsilon_{e\tau}$ and $\epsilon_{\mu\tau}$. Upper panel for the appearance channel and lower panel for the disappearance channel, with no background contributions included. In each panel, the red curve corresponds to the standard oscillation scenario, the green curve to the case with a non-zero real part of the NSI parameter, and the blue curve to the full complex parameter with its associated phase. The shaded regions denote the neutrino flux distribution.
A clear correspondence with the probability level plots of Fig. 1 and Fig. 4 is observed. For the appearance channel ($P_{\mu e}$), deviations in the event rate in Fig. 6 closely follow those seen in the probability plots in Fig. 1. The suppression or enhancement of the blue and green curves relative to the red standard oscillation case is mirrored in the number of electron neutrino events. Around the flux peak ($\sim$2.5 GeV), these deviations are most pronounced.
For the disappearance channel, the event rate spectra shown in the lower panels of Fig. 6 mirror the behavior of the survival probability $P_{\mu\mu}$ in Fig. 4. As in the disappearance probability analysis, the event distributions are largely insensitive to $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$, with the corresponding curves almost overlapping with the standard oscillation case. In contrast, the parameter $\epsilon_{\mu\tau}$ produces a visible shift. Thus, the disappearance event rates clearly follow the same trend as Fig. 4.
Overall, the event-level results confirm that the modifications due to NSI parameters are not independent from those seen in the oscillation probabilities. Instead, the event rates naturally follow the same trends, with the DUNE probability depending on the energy range. Thus, the analysis at the probability level provides the fundamental explanation for the event rate patterns observed in the simulations. For this reason, we only present the event rates to demonstrate their similarity with the corresponding probabilities. We have not performed any sensitivity study or incorporated systematic effects such as flux normalization, cross-section, or detector efficiency uncertainties. Since the main emphasis of this paper is on probabilities rather than on experimental sensitivity, systematic uncertainties are also not considered here.
V.3 Analysis based on quantum entanglement measures
In this subsection, we discuss how the off-diagonal NSI parameters influence the entanglement measures: Entanglement of Formation (EOF), Concurrence, and Negativity. Since all three entanglement measures are functions of neutrino oscillation and survival probabilities, the impact of NSI parameters on these quantities is essentially determined by their influence on the oscillation probabilities. The event rates and the corresponding oscillation probabilities are correlated, as shown in Eq. (30), the overall structure and trends of the event rates follow those of the underlying probabilities. The quantum correlation measures considered in this work cannot be extracted directly from event rates but are expressed solely as functions of these oscillation probabilities. These measures are defined in terms of the neutrino density matrix, which itself depends on the oscillation probabilities. Since the probabilities can be inferred from the measured event rates, the corresponding quantum correlation measures can be accessed indirectly through experimental data, even though they are not directly observable quantities.
Quantum correlation measures have been widely investigated in neutrino systems, consistently revealing nonclassical features in oscillation phenomena. A key motivation behind such studies is to examine whether neutrino oscillations being inherently quantum mechanical could be mimicked or altered by alternative formulations of quantum theory. In this work, we focus on the role of non-standard interactions (NSI) and explore how they affect various quantum correlation measures, with particular attention to entanglement.
The quantities plotted in our work are standard entanglement measures, Entanglement of Formation, Concurrence and Negativity constructed from the neutrino oscillation probabilities. By definition these measures are zero only when the underlying neutrino state is unentangled and become larger as quantum correlations increase; hence the value of the plotted function directly indicates the degree of entanglement between flavor states.
Entanglement provides a complementary lens for studying quantum correlations in neutrino oscillations, as it can be expressed directly in terms of oscillation probabilities. Since our study emphasizes the impact of NSI on entanglement, any modifications to oscillation probabilities induced by NSI are reflected in the entanglement measure. Therefore, identifying NSI effects through probability measurements can be complemented by observing their imprint on entanglement. This dual perspective enriches our understanding of NSI and improves our ability to assess their impact on neutrino dynamics.